OpenStudy (anonymous):

Choose the equation for the hyperbola centered at the origin with the given characteristics. 1.one focus (0, square root of 34), one vertex (0, 5) 2.vertical transverse axis, b = 6, c = square root of 45 3.vertices (+-2, 0), perimeter of central rectangle 24 units

OpenStudy (anonymous):

I have multiple choice selections for these problems

OpenStudy (phi):

The equation of a hyperbola that opens up/down (north/south) is \[ \frac{y^2}{a^2}- \frac{x^2}{b^2} =1 \] I remember that the y term goes first with "Y the smile? Y the frown?" For 1.one focus (0, square root of 34), one vertex (0, 5) we can plot these points  so we know we are looking for \[ \frac{y^2}{a^2}- \frac{x^2}{b^2} =1 \]